Hey there can anyone solve this question?
3X+2Y=12; FIND X WHEN Y=1.5

Answer by jim_thompson5910(35256)

Hey there can anyone solve this question?
3X+2Y=12; FIND X WHEN Y=1.5


3x+2y=12


3x+2(1.5)=12


3x+3=12


3x=12-3


3x=9


x=9/3


x = 3